# Neuro-Vision to Language: Image Reconstruction and Language enabled Interaction via Brain Recordings

### v2

@@ Abstract
Decoding non-invasive brain recordings is crucial for advancing our understanding of human cognition, yet faces challenges from individual differences and complex neural signal representations. Traditional methods require custom models and extensive trials, and lack interpretability in visual reconstruction tasks. Our framework integrating integrates 3D brain structures with visual semantics by Vision Transformer 3D. The unified feature extractor aligns fMRI features with multiple levels of visual embeddings efficiently, removing the need for individual-specific models and allowing extraction from single-trial data. This extractor consolidates multi-level visual features into one network, simplifying integration with Large Language Models (LLMs). Additionally, we have enhanced the fMRI dataset with various fMRI-image related textual data to support multimodal large model development. The integration with LLMs enhances decoding capabilities, enabling tasks like brain captioning, question-answering, detailed descriptions, complex reasoning, and visual reconstruction. Our approach not only shows superior performance across these tasks but also precisely identifies and manipulates language-based concepts within brain signals, enhancing interpretability and providing deeper neural process insights. These advances significantly broaden non-invasive brain decoding applicability in neuroscience and human-computer interaction, setting the stage for advanced brain-computer interfaces and cognitive models.
@@ 1 Introduction
Traditionally, brain decoding techniques have relied on customized, subject-specific models [7; 6; 8]. These models necessitate intricate and costly experimental setups, with a dependency on multiple trials to achieve reliable results. Such approaches, while useful, are inherently limited in scalability and flexibility, hindering broader application and generalization across different populations and conditions.
Visual reconstruction [9] aims to recreate perceived visual stimuli from brain signals and are considered one of the benchmark of brain decoding. However, this approach often struggles to accurately reproduce the visual experiences of individuals, generally lacking semantic precision and interpretability [8]. This inability to effectively decode and reconstruct signals restricts our understanding of how sensory information is processed. Recognizing these limitations, our study introduces language modalities as a key enhancement, designed to assess decoding performance more effectively and to enrich the interaction capabilities of brain-computer interfaces.
@@
Our fMRI feature extractor includes a single unified network backbone along with two alignment heads for feature matching. This setup enables efficient and high-quality visual reconstructions across different subjects from just one experimental trial. By simply aligning the extractor’s output with CLIP embeddings [14] and features of Variational Autoencoders (VAE) [15], our method eliminates the need for multiple, subject-specific models, substantially simplifying the decoding process. This straightforward and effective configuration not only reduces the resources required for brain decoding but also showcases the potential for easy integration with Large Language Models (LLMs), enhancing its usability across various applications.
Moreover, our research delves into the integration of brain recordings with visual and linguistic data within a comprehensive multimodal framework using LLMs. This integration not only significantly improves the performance of visual reconstructions but also introduces the groundbreaking capability for direct interaction through natural language. Our model not only facilitates diverse communication with brain data using natural language but also precisely localizes and manipulates linguistic concepts within brain signals. These advancements help deepen our understanding of the interactions between language, perception, and neural activity. Additionally, to bolster the development of these multimodal models, we have augmented the brain-recording visual dataset with natural language enhancements.
@@
  • Our fMRI feature extractor, based on Vision Transformer 3D, aligns fMRI features with visual embeddings at multiple levels, integrating 3D brain structures with visual semantics. This eliminates the need for individual-specific models and enables efficient data extraction from single trials, significantly reducing training costs and enhancing practical usability in real-world scenarios.
  • We expanded the language dimension of our fMRI-visual dataset to build a multimodal large model capable of decoding fMRI data. This enhancement boosts brain decoding performance and broadens the application scope to include tasks like visual reconstruction, question-answering, and complex reasoning, while also allowing precise localization and manipulation of language-based concepts within brain signals.
  • Experimental results on Natural Scenes Dataset (NSD) [16] for visual reconstruction and language interaction tasks demonstrate that our method surpasses existing models, effectively achieving concept localization and elimination.
@@ 2 Related Works
Non-invasive techniques such as functional magnetic resonance imaging (fMRI) are pivotal in providing direct insights into neural activities, significantly deepening our understanding of complex cognitive processes from neural network structures [17] to advanced image and language processing tasks [18; 19]. This section reviews key developments in fMRI-based brain decoding, particularly emphasizing the shift from simple single-subject analyses to more complex, multimodal interpretations.
@@
Visual Reconstruction from Brain Activity Visual reconstruction from brain activity involves translating brain recordings into the visual stimuli perceived by subjects. Early methods, like those developed by Horikawa et al. [18], relied on sparse linear regression to predict features extracted by convolutional neural networks from fMRI data. Recent advancements in generative artificial intelligence, particularly diffusion models [20], have propelled efforts to reconstruct visual stimuli directly from fMRI. For instance, Lin et al. [21] aligned fMRI data with image features and corresponding CLIP embeddings to facilitate image generation using fine-tuned StyleGAN. Similarly, Takagi et al. [9] improved the quality of visual reconstructions by aligning fMRI with CLIP text embeddings and the latent spaces of diffusion models. Xia et al. [11] aligned fMRI data from dimensions of image CLIP features, depth, and color using a T2I Adapter [22] for fine-grained conditional control. Despite these advancements, the complexity of such methods, involving multiple independent modules, complicates their integration with technologies like LLMs and restricts their generalizability across different subjects.
@@
fMRI Data Processing Efficiently processing fMRI data to extract visually relevant activities typically involves simplifying the data into one-dimensional vectors and selecting voxels most responsive to visual stimuli. Traditional methods utilize simple linear regression or fully connected networks to predict visual stimulus features [18; 21]. However, these methods often lose essential spatial structural information, which is critical given the individual differences in brain anatomy. To address these challenges, innovations such as the Vision Transformer 3D (ViT3D) have been developed for managing data with intricate spatial structures. ViT3D segments 3D data into chunks, preserving local spatial information within each patch and maintaining overall structural integrity through the Transformer’s self-attention mechanism, thereby enhancing the accuracy of brain activity extraction [23].
@@
Multimodal Integration with Brain Signals The utilization of language as a medium for representation allows for the expression of complex concepts with precision and abstraction. The advent of LLMs has showcased their potential to act as bridges across different modalities, enhancing interactions with visual and audio data through natural language [24; 25]. For example, approaches like those by Défossez et al. [26], which align brain recordings with spoken language to decode speech non-invasively, have demonstrated the effectiveness of combining brain recordings with LLMs. However, these approaches are often limited by the specificity of the fMRI feature extractors used, which can restrict scalability and the size of the models employed. By integrating our specially designed cross-subject fMRI feature extractor, we enhance visual reconstruction quality and enable complex reasoning and direct interaction with model outputs using natural language, achieving precise localization of open natural language concepts within the brain.
@@ 3 Methodology
In the sections, we detail the components of our methodology, as seen in Fig. 1. We describe the architecture of our feature preprocessor hat maintains the spatial structure of fMRI data, our unified fMRI feature extractor, and the integration strategies for the network with LLMs. We also elaborate on the multimodal interaction techniques that enable direct and meaningful communication between the computational model and the neural representations, as well as the implementation details for visual reconstruction.
@@
$\left\{p:b_{\text{origin}}^{s}\mapsto b\mid b_{\text{origin}}^{s}\in\mathbb{R}^{X_{s}\times Y_{s}\times Z_{s}},b\in\mathbb{R}^{N\times C}\right\}$ (1)
In Eq. 1, $p$ is the preprocessing function that applies resizing, patching, and masking to transform the original fMRI data into a structured format. This function provides a uniform representation of the BOLD signals across different subjects, ensuring that the spatial structure is not only preserved but also capable of integration with Transformer architectures.
@@ 3.2 Dual-Stream fMRI Feature Extractor
Visual reconstruction tasks typically begin by mapping the processed BOLD signal, $b$, to various estimated visual feature spaces, represented as $\{\hat{z_{1}},\hat{z_{2}},\ldots\}$, where $z$ indicates different levels of visual features and $\hat{z}$ denotes the visual features estimated from the BOLD signals. Subsequently, these features are used to reconstruct the image, represented as $\hat{x}$, from the visual features. To enhance the quality of the reconstructed images, complex feature extractor designs are required, which increases the preprocessing and computational overhead. However, thanks to the design of our fMRI feature extractor, we can achieve efficient visual reconstruction using a single network backbone, as shown in Fig. 1b.
@@
$\mathcal{L}_{\text{align}}=\mathbb{E}_{(b,x)\sim P(B,X)}\left[\left\|f_{w_{c}}(\hat{h}^{N_{b}}_{0})-\mathcal{E}_{c}(x)\right\|_{2}^{2}+\alpha\left\|f_{w_{v}}(\hat{h}^{N_{b}}_{0})-\mathcal{E}_{v}(x)\right\|_{2}^{2}\right]$ (2)
In Eq. 2, the expectation $(b,x)\sim P(B,X)$ averages the alignment loss across samples from the $\bf{B}$ (fMRI signals) and $\bf{X}$ (corresponding visual stimuli). Here, $\hat{h}^{N_{b}}_{0}$ represents the output from the first token of the last hidden state layer of the encoder $\mathcal{E}_{b}$. The functions $f_{w_{c}}$ and $f_{w_{v}}$ are two-layer perceptrons designed to align the extracted fMRI features with these embeddings. The hyperparameter $\alpha$ balances the losses between the alignments of CLIP and VAE features. Through this dual-stream configuration, we create a backbone network that incorporates different levels of visual features. Using only a simple MSE loss function, we achieve high-quality visual reconstruction.
3.3 Multimodal Interaction
Our feature extractor architecture, equipped with a single backbone network, is adept at encapsulating various feature levels, making it highly suitable for integration with LLMs. Inspired by advancements such as those in LLaVA [27], we utilize the penultimate hidden state of our network, $\hat{h}^{N_{b}-1}$, as a multimodal token of fMRI data. A two-layer perceptron $f_{w_{t}}$ projects this state to the same dimension as the text embeddings, resulting in the fMRI embeddings $\hat{z}_{t}=f_{w_{t}}(\hat{h}^{N_{b}-1})$. Considering a sequence of question-answer pairs related to the fMRI data $[q_{0},a_{0},q_{1},a_{1},\ldots,q_{L},a_{L}]$, the training objective is formulated as:
$\max p_{\mathbf{\theta}}(\mathbf{A}|\mathbf{Q},\hat{z}_{t})=\max\prod_{j=0}^{L}p_{\mathbf{\theta}}(a_{j}|q_{j},a_{j-1},\ldots,q_{0},\hat{z}_{t})$ (3)
Eq. 3 describes the probability of generating a sequence of answers $\mathbf{A}$ given a sequence of questions $\mathbf{Q}$ and the derived fMRI embeddings $\hat{z}_{t}$. Each answer $a_{j}$ is conditionally dependent on all preceding questions and answers, as well as the context embeddings derived from fMRI data, where $\theta$ represents the trainable parameters of LLMs.
To effectively couple fMRI data with language models, annotated data is essential. Although the Natural Scenes Dataset (NSD) [16], which uses annotated COCO data, provides labeled data, semantic mismatches often occur due to modifications such as image cropping performed when displaying images to subjects (Appendix A.1). Moreover, the original captions in NSD are not detailed to capture nuanced semantic information. Recognizing the importance of comprehensive linguistic annotations, we have constructed a diverse instructional dataset that includes a variety of textual data: brief descriptions, detailed descriptions, continuous dialogues, and complex reasoning tasks.
@@ 3.4 Interaction and Reconstruction via LLMs
Our fine-tuned model is capable of understanding information embedded within fMRI data and adhering to human instructions. Interactions and explanations of visual stimuli content occur through natural language. A typical dialogue format is structured as follows: <human>:[image] [instruction] <bot>:[answer]. Here, [instruction] denotes a natural language instruction, which during inference is tokenized and embedded, while [image] acts as a placeholder, replaced by the fMRI data embedding $\hat{z}_{t}$. The model then responds based on the directive and the embedded fMRI data. [answer] represents the response generated by the LLMs.
After instruction-based fine-tuning, the model not only communicates directly via natural language but also supports image reconstruction and location identification of concepts expressed in natural language, as shown in Fig. 3. These are facilitated respectively through Stable UnCLIP [20] for image reconstruction and GradCAM [28] for concept localization.
For image reconstruction, the LLM initially generates a simple description of the image $t$, which, combined with the latent representations $z_{v}$ and $z_{c}$ from the fMRI feature extractor, results in the generation of an image. This process can be formalized as:
$\hat{x}=\mathcal{D}((1-\beta)\hat{z}_{v}+\beta\sigma\mid\hat{z}_{c},t),\quad\sigma\in\mathcal{N}(0,1)$ (4)
In Eq. 4, $\mathcal{D}$ represents the frozen UnCLIP [20] model, used for image reconstruction, where $z_{c}$ and $t$ serve as conditional information during the denoising process, and $\hat{z}_{v}$ acts as the initial latent representation of the image. The hyperparameter $\beta$ is used to introduce noise into the latent space prior, balancing low-level features brought by the prior and high-level information controlled by the diffusion conditions during the denoising process. For concept localization in natural language, LLMs activate the feature extractor using keywords from the instructions, enabling precise location identification of the discussed concepts.

### v4

# Neuro-Vision to Language: Enhancing Brain Recording-based Visual Reconstruction and Language Interaction

@@ Abstract
Decoding non-invasive brain recordings is pivotal for advancing our understanding of human cognition but faces challenges due to individual differences and complex neural signal representations. Traditional methods often require customized models and extensive trials, lacking interpretability in visual reconstruction tasks. Our framework integrates 3D brain structures with visual semantics using a Vision Transformer 3D. This unified feature extractor efficiently aligns fMRI features with multiple levels of visual embeddings, eliminating the need for subject-specific models and allowing extraction from single-trial data. The extractor consolidates multi-level visual features into one network, simplifying integration with Large Language Models (LLMs). Additionally, we have enhanced the fMRI dataset with diverse fMRI-image-related textual data to support multimodal large model development. Integrating with LLMs enhances decoding capabilities, enabling tasks such as brain captioning, complex reasoning, concept localization, and visual reconstruction. Our approach demonstrates superior performance across these tasks, precisely identifying language-based concepts within brain signals, enhancing interpretability, and providing deeper insights into neural processes. These advances significantly broaden the applicability of non-invasive brain decoding in neuroscience and human-computer interaction, setting the stage for advanced brain-computer interfaces and cognitive models.
@@ 1 Introduction
Brain decoding techniques have traditionally relied on customized, subject-specific models [7; 6; 8]. These models necessitate intricate and costly experimental setups, depending on multiple trials to achieve reliable results. Such approaches, while helpful, are inherently limited in scalability and flexibility, hindering broader application and generalization across different populations and conditions.
Visual reconstruction [9] aims to recreate perceived visual stimuli from brain signals and is considered one of the benchmarks of brain decoding. However, this approach often struggles to accurately reproduce the visual experiences of individuals, generally lacking semantic precision and interpretability [8]. This inability to effectively decode and reconstruct signals restricts our understanding of how sensory information is processed. Recognizing these limitations, our study introduces language modalities as a critical enhancement designed to assess decoding performance more effectively and enrich brain-computer interfaces’ interaction capabilities.
@@
Our fMRI feature extractor includes a unified network backbone and two alignment heads for feature matching. This setup enables efficient, high-quality visual reconstructions across subjects from one experimental trial. By simply aligning the extractor’s output with CLIP embeddings [14] and features of Variational Autoencoder (VAE) [15], our method eliminates the need for multiple, subject-specific models, substantially simplifying the decoding process. This straightforward and effective configuration reduces the resources required for brain decoding and showcases the potential for easy integration with Large Language Models (LLMs), enhancing its usability across various applications.
Moreover, our research delves into the integration of brain recordings with visual and linguistic data within a comprehensive multimodal framework using LLMs. This integration significantly improves visual reconstruction performance and introduces the groundbreaking capability for direct interaction through natural language. Our model facilitates diverse communication with brain data using natural language and precisely localizes linguistic concepts within brain signals. These advancements help deepen our understanding of the interactions between language, perception, and neural activity. Additionally, to bolster the development of these multimodal models, we have augmented the brain-recording visual dataset with natural language enhancements.
@@
  • Our fMRI feature extractor, based on Vision Transformer 3D, aligns fMRI features with visual embeddings at multiple levels, integrating 3D brain structures with visual semantics. This eliminates the need for subject-specific models and enables efficient data extraction from single trials, significantly reducing training costs and enhancing practical usability in real-world scenarios.
  • We expanded the language dimension of our fMRI-visual dataset to build a multimodal large model capable of decoding fMRI data. This enhancement boosts brain decoding performance and broadens the application scope to include tasks like visual reconstruction, question-answering, and complex reasoning while also allowing precise localization and manipulation of language-based concepts within brain signals.
  • Experimental results on the Natural Scenes Dataset (NSD) [16] for visual reconstruction and language interaction tasks demonstrate that our method surpasses existing models, effectively achieving concept localization and elimination.
@@ 2 Related Works
Non-invasive techniques such as functional magnetic resonance imaging (fMRI) are pivotal in providing direct insights into neural activities, significantly deepening our understanding of complex cognitive processes from neural network structures [17] to advanced image and language processing tasks [18; 19]. This section reviews key developments in fMRI-based brain decoding, particularly emphasizing the shift from simple subject-specific analyses to more complex, multimodal interpretations.
@@
Visual Reconstruction from Brain Activity Visual reconstruction from brain activity involves translating brain recordings into the visual stimuli perceived by subjects. Early methods, like those developed by Horikawa et al. [18], relied on sparse linear regression to predict features extracted by convolutional neural networks from fMRI data. Recent advancements in generative artificial intelligence, particularly diffusion models [20], have propelled efforts to reconstruct visual stimuli directly from fMRI. For instance, Lin et al. [21] aligned fMRI data with image features and corresponding CLIP embeddings to facilitate image generation using fine-tuned StyleGAN [22]. Similarly, Takagi et al. [9] improved the quality of visual reconstructions by aligning fMRI with CLIP text embeddings and the latent spaces of diffusion models. Xia et al. [11] aligned fMRI data from dimensions of image CLIP features, depth, and color using T2I Adapters [23] for fine-grained conditional control. Despite these advancements, the complexity of such methods, involving multiple independent modules, complicates their integration with technologies like LLMs and restricts their generalizability across different subjects. We observed that some contemporary works also attempt cross-subject alignment; however, these methods either require subject-specific parameters [24] or face performance issues compared to subject-specific models [25].
@@
fMRI Data Processing Efficiently processing fMRI data to extract visually relevant activities typically involves simplifying the data into one-dimensional vectors and selecting voxels most responsive to visual stimuli. Traditional methods utilize simple linear regression or fully connected networks to predict visual stimulus features [18; 21]. However, these methods often lose essential spatial structural information, which is critical given the individual differences in brain anatomy. To address these challenges, innovations such as Vision Transformer 3D (ViT3D) have been developed for managing data with intricate spatial structures [26; 27]. ViT3D segments 3D data into patches, preserving local spatial information within each patch and maintaining overall structural integrity through self-attention mechanism [28], thereby enhancing the performance of brain activity extraction [29].
@@
Multimodal Integration with Brain Signals The utilization of language as a medium for representation allows for the expression of complex concepts with precision and abstraction. The advent of LLMs has showcased their potential to act as bridges across different modalities, enhancing interactions with visual and audio data through natural language [30; 31]. For example, approaches like those by Défossez et al. [32], which align brain recordings with spoken language to decode speech non-invasively, have demonstrated the effectiveness of combining brain recordings with LLMs. However, these approaches are often limited by the specificity of the fMRI feature extractors used, which can restrict the scalability and the size of the models employed. By integrating our specially designed cross-subject fMRI feature extractor, we enhance visual reconstruction quality and enable complex reasoning and direct interaction with model outputs using natural language, achieving precise localization of open natural language concepts within the brain.
@@ 3 Methodology
In the following sections, we detail our methodology’s components, as seen in Fig. 1. We describe the architecture of our feature preprocessor that maintains the spatial structure of fMRI data, our unified fMRI feature extractor, and the integration strategies for the network with LLMs. We also elaborate on the multimodal interaction techniques that enable direct and meaningful communication between the computational model and the neural representations and the implementation details for visual reconstruction.
@@
$\left\{p:b_{\text{origin}}^{s}\mapsto b\mid b_{\text{origin}}^{s}\in\mathbb{R}^{X_{s}\times Y_{s}\times Z_{s}},b\in\mathbb{R}^{N\times C}\right\}.$ (1)
In Eq. 1, $p$ is the preprocessing function applies resizing, patching, and masking to transform the original fMRI data into a structured format. This function provides a uniform representation of the BOLD signals across different subjects, ensuring that the spatial structure is preserved and capable of integration with Transformer architectures.
@@ 3.2 Dual-Stream fMRI Feature Extractor
Visual reconstruction tasks typically begin by mapping the processed BOLD signal, $b$, to various estimated visual feature spaces, represented as $\{\hat{z_{1}},\hat{z_{2}},\ldots\}$, where $z$ indicates different levels of visual features and $\hat{z}$ denotes the visual features estimated from the BOLD signals. Subsequently, these features are used to reconstruct the image, represented as $\hat{x}$, from the visual features. To enhance the quality of the reconstructed images, complex feature extractor designs are required, which increases the preprocessing and computational overhead. However, thanks to the design of our fMRI feature extractor, we can achieve efficient visual reconstruction using a single network backbone, as shown in Fig. 1(b).
@@
$\mathcal{L}_{\text{align}}=\mathbb{E}_{(b,x)\sim P(B,X)}\left[\left\|f_{w_{c}}({h}^{N_{b}}_{0})-\mathcal{E}_{c}(x)\right\|_{2}^{2}+\alpha\left\|f_{w_{v}}({h}^{N_{b}}_{0})-\mathcal{E}_{v}(x)\right\|_{2}^{2}\right].$ (2)
In Eq. 2, the expectation $(b,x)\sim P(B,X)$ averages the alignment loss across samples from the $\bf{B}$ (fMRI signals) and $\bf{X}$ (corresponding visual stimuli). Here, ${h}^{N_{b}}_{0}$ represents the output from the first token of the last hidden state layer of the encoder $\mathcal{E}_{b}$. The functions $f_{w_{c}}$ and $f_{w_{v}}$ are two-layer perceptrons designed to align the extracted fMRI features with these embeddings. The hyperparameter $\alpha$ balances the losses between the alignments of CLIP and VAE features. Through this dual-stream configuration, we create a backbone network that incorporates different levels of visual features. Using only a simple MSE loss function, we achieve high-quality visual reconstruction.
3.3 Multimodal Interaction with fMRI
Our feature extractor architecture, equipped with a single backbone network, is adept at encapsulating various feature levels, making it highly suitable for integration with LLMs. Inspired by advancements such as those in LLaVA [33], we utilize the penultimate hidden states of our network, $h^{N_{b}-1}$, as multimodal tokens of fMRI data. A two-layer perceptron $f_{t}$ projects this state to the same dimension as the text embeddings, resulting in the fMRI embeddings $t=f_{t}(h^{N_{b}-1})$. Considering a sequence of question-answer pairs related to the fMRI data $[q_{0},a_{0},q_{1},a_{1},\ldots,q_{L},a_{L}]$, the training objective is formulated as:
$\max p_{\mathbf{\theta}}(\mathbf{A}|\mathbf{Q},t)=\max\prod_{j=0}^{L}p_{\mathbf{\theta}}(a_{j}|q_{j},a_{j-1},\ldots,q_{0},t).$ (3)
Eq. 3 describes the probability of generating a sequence of answers $\mathbf{A}$ given a sequence of questions $\mathbf{Q}$ and the derived fMRI embeddings $t$. Each answer $a_{j}$ is conditionally dependent on all preceding questions and answers, as well as the context embeddings derived from fMRI data, where $\theta$ represents the trainable parameters of LLMs.
To effectively couple fMRI data with language models, annotated data is essential. Although the NSD [16] uses labeled visual stimuli from the COCO dataset [34], semantic mismatches often occur due to modifications such as image cropping performed when displaying images to subjects (Appendix A.1). Moreover, the original captions in NSD are not detailed enough to capture nuanced semantic information. Recognizing the importance of comprehensive linguistic annotations, we have constructed a diverse instructional dataset that includes various textual data: brief descriptions, detailed descriptions, continuous dialogues, complex reasoning tasks, instruction reconstruction, and concept localization.
@@ 3.4 Interaction and Reconstruction via LLMs
Our fine-tuned model can understand information embedded within fMRI data and adhere to human instructions. Interactions and explanations of visual stimuli content occur through natural language. A typical dialogue format is structured as follows: <human>:[image] [instruction] <bot>:[answer]. Here, [instruction] denotes a natural language instruction, which during inference is tokenized and embedded, while [image] acts as a placeholder, replaced by the fMRI data embedding $t$. The model then responds based on the directive and the embedded fMRI data. [answer] represents the response generated by the LLMs.
After instruction-based fine-tuning, the model communicates directly via natural language and supports visual reconstruction and location identification of concepts expressed in natural language, as shown in Fig. 3. These are facilitated respectively through Stable UnCLIP [20] for visual reconstruction and GradCAM [35] for concept localization.
For visual reconstruction, the LLM initially generates a reconstruction prompt $a_{r}$, which, combined with the latent representations $\hat{z}_{v}$ and $\hat{z}_{c}$ from the fMRI feature extractor, results in the generation of an image. This process can be formalized as:
$\hat{x}=\mathcal{D}((1-\beta)\hat{z}_{v}+\beta\sigma\mid\hat{z}_{c},a_{r}),\quad\sigma\in\mathcal{N}(0,1).$ (4)
In Eq. 4, $\mathcal{D}$ represents the frozen UnCLIP, used for visual reconstruction, where $\hat{z}_{c}$ and $a_{r}$ serve as conditional information during the denoising process, and $\hat{z}_{v}$ acts as the initial latent representation of the image. The hyperparameter $\beta$ is used to introduce noise into the latent space prior, balancing low-level features brought by the prior and high-level information controlled by the diffusion conditions during the denoising process. For concept localization in natural language, LLMs activate the feature extractor using keywords from the instructions, enabling precise location identification of the discussed concepts.